If 6 is added to a number and then subtracted from the same nunber, the product of the sum and the difference 189. What is the number? Totally stumped
Answer by jojo14344(1513)

Okay,let's assign "x" as the number.
First theory, the "x" number is added to 6, so (x+6) right?
Second theory, 6 is subtracted from the same "x" number, so (x-6) right?
And the third theory is that the product of the 1st & 2nd theory (sum & diff) is 189.
Putting that into equation:
(x+6)(x-6)=189
x^2-36=189
x^2=189+36
x^2=225
x=15
To check:
(15+6)(15-6)=189
(21)(9)=189
189=189